OpenStudy (ajanijones):

Find the experimental probability. Sheila drew single marbles from a bag, recorded the outcome, and replaced the marble. In 50 draws, Sheila drew 20 white marbles and 30 black marbles. What is the experimental probability that the next marble will be black?

OpenStudy (ajanijones):

A.= 2/5 B. = 3/5 C.2/3 D. 3/7

OpenStudy (ajanijones):

Please help

OpenStudy (anonymous):

number of black balls / total number of balls

OpenStudy (ajanijones):

So do you know the answer cause i dont :(

OpenStudy (ajanijones):

@sourwing do you know the answer please.